# Improved quantum data analysis

### 12 Dec 2023 v3

@@ 1 Introduction
These kinds of questions become much more difficult to analyze when the classical source of randomness $p$ is replaced by a quantum source of randomness, namely a $d$-dimensional quantum state $\rho\in\mathbb{C}^{d\times d}$ (satisfying $\rho\geq 0$, $\operatorname{tr}(\rho)=1$). The difficulties here are that: (i) one cannot directly observe “outcomes” for $\rho$, one can only measure it; (ii) measuring the state $\rho$ inherently alters it, hence reusing samples (i.e., copies of $\rho$) is problematic. For example, suppose we now have some known, fixed observables $A_{1},\dots,A_{m}\in\mathbb{C}^{d\times d}$ with $0\leq A_{i}\leq{\mathbbm{1}}$ and we wish to estimate each expectation $\operatorname*{\mathbf{E}}_{\rho}[A_{i}]\coloneqq\operatorname{tr}(\rho A_{i})$ to within ±$\pm\epsilon$. This is the “Shadow Tomography” problem introduced by Aaronson in [Aar16] (see [Aar20] for applications to, e.g., quantum money). We do not know if this is similarly possible using $n=O((\log m)/\epsilon^{2})$ copies of $\rho$; indeed, prior to this work the best known upper bound was
@@
Here the sample complexity on the left is from [Aar20], combining a “Gentle Search” routine with an online learning algorithm for quantum states from [${}^{+}$19]. The sample complexity on the right was obtained by Aaronson and Rothblum [AR19] by drawing inspiration and techniques from the field of Differential Privacy. [^1] See also [HKP20] for sample complexity bounds that can be better for special kinds of $A_{i}$’s.
In fact, we propose that — rather than Differential Privacy — a closer classical match for the Shadow Tomography problem is the task known as Adaptive Data Analysis, introduced by [${}^{+}$15]. In this problem, the random variables (“statistical queries”) $a_{1},\dots,a_{m}$ are not fixed in advance for the learner, but are rather received one at a time, with the crucial feature that each $a_{t}$ may adaptively depend on the preceding estimates of $\operatorname*{\mathbf{E}}_{p}[a_{1}],\dots,\operatorname*{\mathbf{E}}_{p}[a_{t-1}]$ output by the learner. In this case, conditioning on these output estimates skews the underlying i.i.d. product distribution $p^{\otimes n}$ — reminiscent of the way measuring a quantum state affects it — and this prevents naive reuse of the sample data. Indeed, it’s far from obvious that the Adaptive Data Analysis task is doable with $\operatorname{poly}(\log m,\log d,1/\epsilon)$ samples; however this was shown by [${}^{+}$15], who achieved complexity $n=\widetilde{O}((\log m)^{3/2}(\log d)^{1/2}/\epsilon^{7/2})$, and this was later improved by [${}^{+}$21] to $n=\widetilde{O}((\log m)(\log d)^{1/2}/\epsilon^{3}))$. While Differential Privacy tools have been an ingredient in some Adaptive Data Analysis routines, the topics are not inherently linked; e.g., a viewpoint based on “KL-stability” is emphasized in [${}^{+}$21].
@@ 1.1.1 Threshold Search
The first main result in our work concerns what we will call the quantum “Threshold Search” problem. [^2] Originally called the “Secret Acceptor” problem when it was introduced by Aaronson [Aar16]. Later he called it “Gentle Search” [Aar20], but we find this name unsatisfactory as it is not necessary that a successful algorithm be “gentle”. In the Differential Privacy literature, it is sometimes called “Report Noisy Max” (offline case) or “Above Threshold” (online case) [DR13]. We state the problem here in a general form (recalling our notation $\operatorname*{\mathbf{E}}_{\rho}[A_{i}]=\operatorname{tr}(\rho A_{i})$):
@@
We remark that all of the difficulty of the problem is contained in the case where $\epsilon=\delta=\frac{1}{4}$ and $\theta_{j}=\frac{3}{4}$ for all $j$ (see Section 4.1). In this case, Aaronson [Aar16] originally showed that the Threshold Search problem can be solved using $n=\widetilde{O}(\log^{4}m)$ copies of $\rho$. In the present paper, we improve this result quadratically:
@@
Incidentally, the (offline) quantum Threshold Decision problem, where the algorithm only needs to report “$\exists j:\operatorname*{\mathbf{E}}_{\rho}[A_{j}]>\theta_{j}-\epsilon$” without actually specifying $j$, is known to be solvable using just $n=O(\log(m)\log(1/\delta)/\epsilon^{2})$ copies [Aar20]. We review the proof in Appendix A, tightening/simplifying some quantitative aspects of the underlying theorem of Harrow, Lin, and Montanaro [HLM17]. In particular, our tightenings let us slightly improve the copy complexity to $n=O(\log(m/\delta)/\epsilon^{2})$.
@@ 1.1.2 $\chi^{2}$-stable threshold reporting
The most important technical ingredient going into our proof of Theorem 1.1 is a new, purely classical statistical result fitting into the Adaptive Data Analysis framework (see, e.g., [Smi17] for some background). In that setting one might describe our result as follows: “adding exponential noise provides a (composably) $\chi^{2}$-stable mechanism for reporting if a distribution’s mean is above a given threshold”. In more detail, the result says that given a Sample ${\boldsymbol{S}}$ consisting of the sum of $n$ draws from a $\mathrm{Bernoulli}(p)$ distribution (i.e., ${\boldsymbol{S}}\sim\operatorname{Binomial}(n,p)$), if we add independent exponential noise ${\boldsymbol{X}}$ and then check the event $B$ that ${\boldsymbol{S}}+{\boldsymbol{X}}$ exceeds some large threshold $\theta n$, then conditioning on $B$ not occurring hardly changes the distribution of ${\boldsymbol{S}}$, provided $\operatorname*{\mathbf{E}}[{\boldsymbol{X}}]\gg\operatorname*{\mathbf{stddev}}[{\boldsymbol{S}}]$. Here the phrase “hardly changes” is in two very strong senses: (i) we show the random variables ${\boldsymbol{S}}\mid\overline{B}$ and ${\boldsymbol{S}}$ are close even in $\chi^{2}$-divergence, which is a more stringent measure than KL-divergence (or Hellinger distance, or total variation distance) — that is, the test is “$\chi^{2}$-stable”; (ii) the $\chi^{2}$-divergence is not just absolutely small, but is even a small fraction of $\operatorname*{\mathbf{P}}[B]^{2}$ itself (hence the total variation closeness is a small fraction of $\operatorname*{\mathbf{P}}[B]$). This allows a kind of composition (as in the “Sparse Vector” mechanism [DR13] from the Differential Privacy literature) in which the same quantum sample can be reused for repeated “above threshold” tests, up until the point where having at least one “above threshold” outcome becomes likely. Precisely, our result is the following (refer to Section 2.1 for the definition and notation of the quantities used below):
@@
Writing ${\boldsymbol{S}}^{\prime}$ for ${\boldsymbol{S}}\mid\overline{B}$, standard inequalities for $f$-divergences [GS02] imply
@@
Let us remark that our Theorem 1.2 is similar to results appearing previously in the Differential Privacy/Adaptive Data Analysis literature; in particular, it is quite similar to (and inspired by) a theorem (“Claim 41”) of Aaronson and Rothblum [AR19]. Although this Claim 41 is presented in a quantum context, the essence of it is a theorem comparable to our Theorem 1.2, with the following main differences: (i) it bounds the weaker KL-divergence (though for our applications, this is acceptable); (ii) the proof is significantly more involved. (Minor differences include: (i) their result uses two-sided exponential noise for a two-sided threshold event; (ii) our bound has the stronger factor $\operatorname*{\mathbf{stddev}}[{\boldsymbol{S}}]$ instead of just $\sqrt{n}$.)
@@ 1.1.3 Applications: Shadow Tomography and Hypothesis Selection
The proof of this theorem is almost immediate from our Threshold Search algorithm, using a known [Aar20] black-box reduction to the mistake-bounded online quantum state learning algorithm of Aaronson, Chen, Hazan, Kale, and Nayak [${}^{+}$19].
@@
The classical (multiple) Hypothesis Selection problem [Yat85, DL96, DL97] is as follows: Given are $m$ fixed “hypothesis” probability distributions $q_{1},\dots,q_{m}$ on $[d]$, as well as a parameter $\epsilon$ and access to samples from an unknown distribution $p$ on $[d]$. The task is to find (with probability at least $1-\delta$) a $q_{j}$ which is, roughly, closest to $p$, while minimizing the number of samples drawn from $p$. More precisely, if $\eta=\min_{i}\{d_{\mathrm{\mathrm{TV}}}(p,q_{i})\}$, the algorithm should output a hypothesis $q_{j}$ with $d_{\mathrm{\mathrm{TV}}}(p,q_{j})\leq C\eta+\epsilon$ for some fixed small constant $C$. There are a variety of solutions known to this problem, with standard ones [DL01, Chap. 6] achieving $n=O((\log m)/\epsilon^{2})$ (and best constant $C=3$). There are also numerous variations, including handling different distance measures besides $d_{\mathrm{\mathrm{TV}}}$ [BV08], the easier (“realizable/non-robust”) case when $\eta=0$, and the case when there is a unique answer (as when the hypotheses $q_{j}$ are pairwise far apart). We emphasize that our focus is on the non-asymptotic regime, where we would like an explicit sample bound $n=n(m,d,\epsilon,\delta)$ holding for all values of $m,d,\epsilon,\delta$ [^3] This is as opposed to the asymptotic regime. There, one focuses on achieving $\delta\leq\exp(-C(m,d,\epsilon)n)$ for all $n\geq n_{0}(m,d,\epsilon)$, where the rate function $C(m,d,\epsilon)$ should be as large as possible, but where $n_{0}$ may be a completely uncontrolled function of $m,d,\epsilon$. See, e.g., [NS11]. One particularly useful application of Hypothesis Selection is to learning an unknown probability distribution $p$ from a class $\mathcal{C}$ (even “agnostically”). Roughly speaking, if $\mathcal{C}$ has an $\epsilon$-cover of size $m=m(\epsilon)$, then one can learn $p$ to accuracy $O(\epsilon)$ using a Hypothesis Selection over $m$ hypotheses; i.e., with $O((\log m)/\epsilon^{2})$ samples in the classical case. For further discussion of the problem, see e.g. [MŠ07]; for Differentially Private Hypothesis Selection, see [BKSW21, ${}^{+}$20]; for fast classical Hypothesis Selection with a quantum computer, see [QCR20].
@@
There is a quantum algorithm that, given $m$ fixed hypothesis states $\sigma_{1},\dots\sigma_{m}\in\mathbb{C}^{d\times d}$, parameters $0<\epsilon,\delta<\frac{1}{2}$, and access to unentangled copies of a state $\rho\in\mathbb{C}^{d\times d}$, uses
@@
The fact that quantum Hypothesis Selection black-box reduces to Shadow Tomography provides significant motivation for trying to prove (or disprove) that Shadow Tomography can be done with $O(\log m)\cdot\operatorname{poly}((\log d)/\epsilon)$ copies; i.e., that the power on $\log m$ can be reduced to $1$. If this were possible, then as in the classical case we would be able to learn a quantum state $\rho\in\mathbb{C}^{d\times d}$ in a class $\mathcal{C}$ (to constant trace distance accuracy, say) using $\log(|\mathrm{cover}(\mathcal{C})|)\cdot\operatorname{polylog}(d)$ copies, where $\mathrm{cover}(\mathcal{C})$ denotes a set of states that form a (trace-distance) cover for $\mathcal{C}$. It’s easy to see that the class $\mathcal{C}$ of all states has a cover of size at most $O(d)^{d^{2}}$, and hence Shadow Tomography with a $\log m$ dependence would yield a full quantum tomography algorithm with copy complexity $\widetilde{O}(d^{2})$, bypassing the sophisticated representation-theory methods of [OW16, ${}^{+}$17, OW17]. One might also hope for more efficient learning of other interesting subclasses of states; e.g., the class separable states.
@@ 2.2 Quantum states and measurements
More generally, we have the mathematical notion of a quantum operation $S$ on $d$-dimensional states, defined by $d$-column matrices $M_{1},\dotsc,M_{k}$ such that
@@ 2.3 Quantum state distances
Just as with classical probability distributions, there are a variety of distances/divergences between two quantum states $\rho,\sigma\in\mathbb{C}^{d\times d}$. In fact, for every classical “$f$-divergence” there is a corresponding “measured quantum $f$-divergence”, which is the maximal classical divergence that can be achieved by performing the same measurement on $\rho$ and $\sigma$. In this way, classical total variation distance precisely corresponds to quantum trace distance, the Bhattacharyya coefficient precisely corresponds to quantum fidelity, etc. See, e.g., [BOW19, Sec. 3.1.2] for further review; here we will simply directly define some quantum distances.
@@
Here the second equality is known as the Holevo–Helstrom theorem [Hol73, Hel76], and the maximum is over all quantum events $A\in\mathbb{C}^{d\operatorname{\times}d}$. Moreover, the maximum is achieved by a projector. The fidelity $\operatorname{F}(\rho,\sigma)$ between states $\rho$ and $\sigma$ is defined by
@@
It follows from the work of Fuchs and Caves [FC95] that $\frac{1}{2}d_{\mathrm{Bures^{2}}}(\rho,\sigma)\leq d_{\mathrm{tr}}(\rho,\sigma)\leq d_{\mathrm{Bures}}(\rho,\sigma)$ for all states $\rho$ and $\sigma$.
Below we give a simpler formula for fidelity in the case when $\sigma$ is a conditioned version of $\rho$ (such results are sometimes known under the name “gentle measurement”; see [Wat18, Cor. 3.15]):
@@
$\operatorname{F}(\rho,{\left.\kern-1.2pt{\rho}\right|}_{M})^{2}=\frac{\operatorname{tr}\left\lparen\sqrt{\sqrt{\rho}M\rho M^{\dagger}\sqrt{\rho}}\right\rparen^{2}}{\operatorname*{\mathbf{E}}_{\rho}[M^{\dagger}M]}=\frac{\operatorname{tr}\left\lparen\sqrt{\sqrt{\rho}M\sqrt{\rho}\sqrt{\rho}M\sqrt{\rho}}\right\rparen^{2}}{\operatorname*{\mathbf{E}}_{\rho}[M^{2}]}=\frac{\operatorname{tr}\left\lparen\sqrt{\rho}M\sqrt{\rho}\right\rparen^{2}}{\operatorname*{\mathbf{E}}_{\rho}[M^{2}]}=\frac{\operatorname*{\mathbf{E}}_{\rho}[M]^{2}}{\operatorname*{\mathbf{E}}_{\rho}[M^{2}]}.\qed$
Below we give a further formula for $\operatorname{F}(\rho,{\left.\kern-1.2pt{\rho}\right|}_{\sqrt{A}})$ using the spectral decomposition of $A$. (We remark that it may be obtained as a special case of the theorem of Fuchs and Caves [FC95].)
@@
and the proof is complete by Proposition 2.2. ∎

### 15 Mar 2024 v4

@@ 1 Introduction
These kinds of questions become much more difficult to analyze when the classical source of randomness $p$ is replaced by a quantum source of randomness, namely a $d$-dimensional quantum state $\rho\in\mathbb{C}^{d\times d}$ (satisfying $\rho\geq 0$, $\operatorname{tr}(\rho)=1$). The difficulties here are that: (i) one cannot directly observe “outcomes” for $\rho$, one can only measure it; (ii) measuring the state $\rho$ inherently alters it, hence reusing samples (i.e., copies of $\rho$) is problematic. For example, suppose we now have some known, fixed observables $A_{1},\dots,A_{m}\in\mathbb{C}^{d\times d}$ with $0\leq A_{i}\leq{\mathbbm{1}}$ and we wish to estimate each expectation $\operatorname*{\mathbf{E}}_{\rho}[A_{i}]\coloneqq\operatorname{tr}(\rho A_{i})$ to within ±$\pm\epsilon$. This is the “Shadow Tomography” problem introduced by Aaronson in [2] (see [1] for applications to, e.g., quantum money). We do not know if this is similarly possible using $n=O((\log m)/\epsilon^{2})$ copies of $\rho$; indeed, prior to this work the best known upper bound was
@@
Here the sample complexity on the left is from [1], combining a “Gentle Search” routine with an online learning algorithm for quantum states from [3]. The sample complexity on the right was obtained by Aaronson and Rothblum [4] by drawing inspiration and techniques from the field of Differential Privacy. [^1] See also [28] for sample complexity bounds that can be better for special kinds of $A_{i}$’s.
In fact, we propose that — rather than Differential Privacy — a closer classical match for the Shadow Tomography problem is the task known as Adaptive Data Analysis, introduced by [17]. In this problem, the random variables (“statistical queries”) $a_{1},\dots,a_{m}$ are not fixed in advance for the learner, but are rather received one at a time, with the crucial feature that each $a_{t}$ may adaptively depend on the preceding estimates of $\operatorname*{\mathbf{E}}_{p}[a_{1}],\dots,\operatorname*{\mathbf{E}}_{p}[a_{t-1}]$ output by the learner. In this case, conditioning on these output estimates skews the underlying i.i.d. product distribution $p^{\otimes n}$ — reminiscent of the way measuring a quantum state affects it — and this prevents naive reuse of the sample data. Indeed, it’s far from obvious that the Adaptive Data Analysis task is doable with $\operatorname{poly}(\log m,\log d,1/\epsilon)$ samples; however this was shown by [17], who achieved complexity $n=\widetilde{O}((\log m)^{3/2}(\log d)^{1/2}/\epsilon^{7/2})$, and this was later improved by [7] to $n=\widetilde{O}((\log m)(\log d)^{1/2}/\epsilon^{3}))$. While Differential Privacy tools have been an ingredient in some Adaptive Data Analysis routines, the topics are not inherently linked; e.g., a viewpoint based on “KL-stability” is emphasized in [7].
@@ 1.1.1 Threshold Search
The first main result in our work concerns what we will call the quantum “Threshold Search” problem. [^2] Originally called the “Secret Acceptor” problem when it was introduced by Aaronson [2]. Later he called it “Gentle Search” [1], but we find this name unsatisfactory as it is not necessary that a successful algorithm be “gentle”. In the Differential Privacy literature, it is sometimes called “Report Noisy Max” (offline case) or “Above Threshold” (online case) [18]. We state the problem here in a general form (recalling our notation $\operatorname*{\mathbf{E}}_{\rho}[A_{i}]=\operatorname{tr}(\rho A_{i})$):
@@
We remark that all the difficulty of the problem is contained in the case where $\epsilon=\delta=\frac{1}{4}$ and $\theta_{j}=\frac{3}{4}$ for all $j$ (see Section 4.1). In this case, Aaronson [2] originally showed that the Threshold Search problem can be solved using $n=\widetilde{O}(\log^{4}m)$ copies of $\rho$. In the present paper, we improve this result quadratically:
@@
Incidentally, the (offline) quantum Threshold Decision problem, where the algorithm only needs to report “$\exists j:\operatorname*{\mathbf{E}}_{\rho}[A_{j}]>\theta_{j}-\epsilon$” without actually specifying $j$, is known to be solvable using just $n=O(\log(m)\log(1/\delta)/\epsilon^{2})$ copies [1]. We review the proof in Appendix A, tightening/simplifying some quantitative aspects of the underlying theorem of Harrow, Lin, and Montanaro [24]. In particular, our tightenings let us slightly improve the copy complexity to $n=O(\log(m/\delta)/\epsilon^{2})$.
@@ 1.1.2 $\chi^{2}$-stable threshold reporting
The most important technical ingredient going into our proof of Theorem 1.1 is a new, purely classical statistical result fitting into the Adaptive Data Analysis framework (see, e.g., [38] for some background). In that setting one might describe our result as follows: “adding exponential noise provides a (composably) $\chi^{2}$-stable mechanism for reporting if a distribution’s mean is above a given threshold”. In more detail, the result says that given a Sample ${\boldsymbol{S}}$ consisting of the sum of $n$ draws from a $\mathrm{Bernoulli}(p)$ distribution (i.e., ${\boldsymbol{S}}\sim\operatorname{Binomial}(n,p)$), if we add independent exponential noise ${\boldsymbol{X}}$ and then check the event $B$ that ${\boldsymbol{S}}+{\boldsymbol{X}}$ exceeds some large threshold $\theta n$, then conditioning on $B$ not occurring hardly changes the distribution of ${\boldsymbol{S}}$, provided $\operatorname*{\mathbf{E}}[{\boldsymbol{X}}]\gg\operatorname*{\mathbf{stddev}}[{\boldsymbol{S}}]$. Here the phrase “hardly changes” is in two very strong senses: (i) we show the random variables ${\boldsymbol{S}}\mid\overline{B}$ and ${\boldsymbol{S}}$ are close even in $\chi^{2}$-divergence, which is a more stringent measure than KL-divergence (or Hellinger distance, or total variation distance) — that is, the test is “$\chi^{2}$-stable”; (ii) the $\chi^{2}$-divergence is not just absolutely small, but is even a small fraction of $\operatorname*{\mathbf{P}}[B]^{2}$ itself (hence the total variation closeness is a small fraction of $\operatorname*{\mathbf{P}}[B]$). This allows a kind of composition (as in the “Sparse Vector” mechanism [18] from the Differential Privacy literature) in which the same quantum sample can be reused for repeated “above threshold” tests, up until the point where having at least one “above threshold” outcome becomes likely. Precisely, our result is the following (refer to Section 2.1 for the definition and notation of the quantities used below):
@@
Writing ${\boldsymbol{S}}^{\prime}$ for ${\boldsymbol{S}}\mid\overline{B}$, standard inequalities for $f$-divergences [21] imply
@@
Let us remark that our Theorem 1.2 is similar to results appearing previously in the Differential Privacy/Adaptive Data Analysis literature; in particular, it is quite similar to (and inspired by) a theorem (“Claim 41”) of Aaronson and Rothblum [4]. Although this Claim 41 is presented in a quantum context, the essence of it is a theorem comparable to our Theorem 1.2, with the following main differences: (i) it bounds the weaker KL-divergence (though for our applications, this is acceptable); (ii) the proof is significantly more involved. (Minor differences include: (i) their result uses two-sided exponential noise for a two-sided threshold event; (ii) our bound has the stronger factor $\operatorname*{\mathbf{stddev}}[{\boldsymbol{S}}]$ instead of just $\sqrt{n}$.)
@@ 1.1.3 Applications: Shadow Tomography and Hypothesis Selection
The proof of this theorem is almost immediate from our Threshold Search algorithm, using a known [1] black-box reduction to the mistake-bounded online quantum state learning algorithm of Aaronson, Chen, Hazan, Kale, and Nayak [3].
@@
The classical (multiple) Hypothesis Selection problem [40, 13, 15] is as follows: Given are $m$ fixed “hypothesis” probability distributions $q_{1},\dots,q_{m}$ on $[d]$, as well as a parameter $\epsilon$ and access to samples from an unknown distribution $p$ on $[d]$. The task is to find (with probability at least $1-\delta$) a $q_{j}$ which is, roughly, closest to $p$, while minimizing the number of samples drawn from $p$. More precisely, if $\eta=\min_{i}\{d_{\mathrm{\mathrm{TV}}}(p,q_{i})\}$, the algorithm should output a hypothesis $q_{j}$ with $d_{\mathrm{\mathrm{TV}}}(p,q_{j})\leq C\eta+\epsilon$ for some fixed small constant $C$. There are a variety of solutions known to this problem, with standard ones [14, Chap. 6] achieving $n=O((\log m)/\epsilon^{2})$ (and best constant $C=3$). There are also numerous variations, including handling different distance measures besides $d_{\mathrm{\mathrm{TV}}}$ [9], the easier (“realizable/non-robust”) case when $\eta=0$, and the case when there is a unique answer (as when the hypotheses $q_{j}$ are pairwise far apart). We emphasize that our focus is on the non-asymptotic regime, where we would like an explicit sample bound $n=n(m,d,\epsilon,\delta)$ holding for all values of $m,d,\epsilon,\delta$ [^3] This is as opposed to the asymptotic regime. There, one focuses on achieving $\delta\leq\exp(-C(m,d,\epsilon)n)$ for all $n\geq n_{0}(m,d,\epsilon)$, where the rate function $C(m,d,\epsilon)$ should be as large as possible, but where $n_{0}$ may be a completely uncontrolled function of $m,d,\epsilon$. See, e.g., [32]. One particularly useful application of Hypothesis Selection is to learning an unknown probability distribution $p$ from a class $\mathcal{C}$ (even “agnostically”). Roughly speaking, if $\mathcal{C}$ has an $\epsilon$-cover of size $m=m(\epsilon)$, then one can learn $p$ to accuracy $O(\epsilon)$ using a Hypothesis Selection over $m$ hypotheses; i.e., with $O((\log m)/\epsilon^{2})$ samples in the classical case. For further discussion of the problem, see e.g. [30]; for Differentially Private Hypothesis Selection, see [11, 22]; for fast classical Hypothesis Selection with a quantum computer, see [37].
@@
There is a quantum algorithm that, given $m$ fixed hypothesis states $\sigma_{1},\dots,\sigma_{m}\in\mathbb{C}^{d\times d}$, parameters $0<\epsilon,\delta<\frac{1}{2}$, and access to unentangled copies of a state $\rho\in\mathbb{C}^{d\times d}$, uses
@@
The fact that quantum Hypothesis Selection black-box reduces to Shadow Tomography provides significant motivation for trying to prove (or disprove) that Shadow Tomography can be done with $O(\log m)\cdot\operatorname{poly}((\log d)/\epsilon)$ copies; i.e., that the power on $\log m$ can be reduced to $1$. If this were possible, then as in the classical case we would be able to learn a quantum state $\rho\in\mathbb{C}^{d\times d}$ in a class $\mathcal{C}$ (to constant trace distance accuracy, say) using $\log(|\mathrm{cover}(\mathcal{C})|)\cdot\operatorname{polylog}(d)$ copies, where $\mathrm{cover}(\mathcal{C})$ denotes a set of states that form a (trace-distance) cover for $\mathcal{C}$. It’s easy to see that the class $\mathcal{C}$ of all states has a cover of size at most $O(d)^{d^{2}}$, and hence Shadow Tomography with a $\log m$ dependence would yield a full quantum tomography algorithm with copy complexity $\widetilde{O}(d^{2})$, bypassing the sophisticated representation-theory methods of [34, 23, 35]. One might also hope for more efficient learning of other interesting subclasses of states; e.g., the class separable states.
@@ 2.2 Quantum states and measurements
More generally, we have the notion of a quantum operation $S$ on $d$-dimensional states, defined by $d$-column matrices $M_{1},\dotsc,M_{k}$ such that
@@ 2.3 Quantum state distances
Just as with classical probability distributions, there are a variety of distances/divergences between two quantum states $\rho,\sigma\in\mathbb{C}^{d\times d}$. In fact, for every classical “$f$-divergence” there is a corresponding “measured quantum $f$-divergence”, which is the maximal classical divergence that can be achieved by performing the same measurement on $\rho$ and $\sigma$. In this way, classical total variation distance precisely corresponds to quantum trace distance, the Bhattacharyya coefficient precisely corresponds to quantum fidelity, etc. See, e.g., [6, Sec. 3.1.2] for further review; here we will simply directly define some quantum distances.
@@
Here the second equality is known as the Holevo–Helstrom theorem [27, 26], and the maximum is over all quantum events $A\in\mathbb{C}^{d\operatorname{\times}d}$. Moreover, the maximum is achieved by a projector. The fidelity $\operatorname{F}(\rho,\sigma)$ between states $\rho$ and $\sigma$ is defined by
@@
It follows from the work of Fuchs and Caves [19] that $\frac{1}{2}d_{\mathrm{Bures^{2}}}(\rho,\sigma)\leq d_{\mathrm{tr}}(\rho,\sigma)\leq d_{\mathrm{Bures}}(\rho,\sigma)$ for all states $\rho$ and $\sigma$.
Below we give a simpler formula for fidelity in the case when $\sigma$ is a conditioned version of $\rho$ (such results are sometimes known under the name “gentle measurement”; see [39, Cor. 3.15]):
@@
$\operatorname{F}(\rho,{\left.\kern-1.2pt{\rho}\right|}_{M})^{2}=\frac{\operatorname{tr}\left\lparen\sqrt{\sqrt{\rho}M\rho M^{\dagger}\sqrt{\rho}}\right\rparen^{2}}{\operatorname*{\mathbf{E}}_{\rho}[M^{\dagger}M]}=\frac{\operatorname{tr}\left\lparen\sqrt{\sqrt{\rho}M\sqrt{\rho}\sqrt{\rho}M\sqrt{\rho}}\right\rparen^{2}}{\operatorname*{\mathbf{E}}_{\rho}[M^{2}]}=\frac{\operatorname{tr}\left\lparen\sqrt{\rho}M\sqrt{\rho}\right\rparen^{2}}{\operatorname*{\mathbf{E}}_{\rho}[M^{2}]}=\frac{\operatorname*{\mathbf{E}}_{\rho}[M]^{2}}{\operatorname*{\mathbf{E}}_{\rho}[M^{2}]}.$
Below we give a further formula for $\operatorname{F}(\rho,{\left.\kern-1.2pt{\rho}\right|}_{\sqrt{A}})$ using the spectral decomposition of $A$. (We remark that it may be obtained as a special case of the theorem of Fuchs and Caves [19].)
@@
and the proof is complete by Proposition 2.2.